# Combining Evidence Across Filtrations

### v4

@@ 4 Experiments: Sequentially testing randomness
Given the $\mathfrak{e}$-lifting result (Theorem 2), we can now revisit the task of sequentially testing randomness, for which the issue of combining e-processes across filtrations arises. In this section, we first revisit our motivating example from Section 1.2 and show results on simulated data. Then, we apply our approach to testing the randomness of real-world financial time series data.
4.1 Simulation results under the null and alternative hypotheses
@@
Recall from Section 1.2 that we could not simply average the two e-processes $\mathfrak{e}^{\mathsf{UI}}$ and $\mathfrak{e}^{\mathsf{conf}}$ for testing randomness, as the conformal test martingale (CTM) $\mathfrak{e}^{\mathsf{conf}}$ is not $\mathbb{F}$-anytime-valid. By Corollary 2, we can now combine the two e-process by lifting $\mathfrak{e}^{\mathsf{conf}}$ into $\mathbb{F}$ using an adjuster.
@@
As shown in the plots, unlike the original CTM, its adjusted version is mostly no greater than 1. Over $10,000$ repeated runs, the mean of the stopped e-values for the adjusted e-process is $0.60\pm 0.001$, suggesting that the adjusted e-process meets the validity definition (1) at this $\mathbb{F}$-stopping time. This gives evidence that the adjusted e-process can be combined with the UI e-process via averaging, such that the combined version is an e-process in $\mathbb{F}$.
@@
Next, we generate data from a drifting distribution that is typically an IID $\mathsf{Ber}(0.5)$ (“usual state”), but after the first 500 steps it briefly switches to an IID $\mathsf{Ber}(0.2)$ (“unusual state”) for 100 steps, before returning to the usual state for the remainder of the sequence. The resulting data contains two changepoints over the sequence of length $T=2,500$. As mentioned earlier, CTMs are known to be powerful against changepoint alternatives, unlike the UI e-process. Figure 5(b), which shows the three e-processes averaged over 1,000 repeated runs, confirms this: the CTM grows large upon entering the unusual state and then gradually decreases upon reentering the usual state. In contrast, the UI e-process remains small throughout the sequence. Because one of the two processes grows large (in the unusual state), analogous to the Markovian case, the combined e-process also grows large.
Unlike in Figure 5(a), however, in Figure 5(b) we additionally observe two peculiarities stemming from the $\mathfrak{e}$-lifting procedure. First, because the combined e-process utilizes the adjusted CTM, it grows less quickly ($\approx 10^{2}$ at $t=500$) than the CTM ($\approx 10^{4}$). In this example, the adjusted evidence still exceeds a typical threshold for a “strong” evidence (Shafer,, 2021). (When using an e-process for a level-$\alpha$ sequential test (7), setting $\alpha$ to $0.1$, $0.05$, or $0.01$ corresponds to rejecting the null when the e-value exceeds 10, 20, or 100, respectively.) Second, as a byproduct of using an adjuster, the combined e-process retains the largest evidence from the CTM over time (the running maximum). As the data returns to its usual state (past time 600), the CTM gradually decreases, while the combined e-process eventually exceeds the CTM. Thus, while adjusting the e-process involves sacrificing some amount of evidence, it also comes with the useful byproduct of being able to insure against losing evidence over time.
@@ 4.2 Testing the randomness of high-volatility days in financial returns
To test the viability of the adjust-then-combine approach in real-world applications, we revisit and expand upon the motivating example from Section 1.2 involving financial time series data. As briefly described earlier, modeling the volatility of returns is a key problem in financial modeling, and a good volatility model is instrumental in tasks such as portfolio management, risk management, and options pricing (Engle and Patton,, 2007). When considering a volatility model, it is useful to accompany it with a monitoring procedure that can check this model’s viability on live data, which may change unpredictably over time.
The problem of sequentially testing whether high-volatility days are distributed at random, as described in Example 1, is a concrete example of this monitoring task. A “stylized fact” about financial returns is that they are heteroskedastic, and specifically, that high-volatility days are often clustered together (e.g., Engle,, 1982). This makes it practically important to detect exactly when (and how) volatility clustering happens on live data. The task can be viewed as a diagnostic test for whether a volatility model is needed in the first place: if high-volatility days are IID, then we would not expect a volatility model to provide added predictive power.
@@
In this experiment, we choose $c$ as the 80% upper quantile of the historical returns for the same asset, from a 1-year period before the first date (say, during the year 2018). In Figure 6, we plot the volatility process $V_{t}$ and the high-volatility days $X_{t}$ (marked as red) for the S&P 500 index, which shows volatility clustering around March 2020 and during 2022.
@@
While we focused on a binarized notion of volatility for simplicity, we can apply an analogous procedure to the continuous-valued residuals from a volatility model, by replacing the UI e-process with Saha and Ramdas, (2024)’s e-process (powerful against autoregressive alternatives) and using the continuous analog of the CTM (Vovk, 2021b, ).
@@ 5.1 Combining e-processes for sequential independence testing
The first example is sequential (nonparametric) independence testing (Balsubramani and Ramdas,, 2016; Podkopaev et al.,, 2023; Podkopaev and Ramdas,, 2023; Shekhar and Ramdas,, 2024; Henzi and Law,, 2024). This represents a scenario where both e-processes are constructed in a coarsening of the data filtration. By a complete analogy to Corollary 2, we can combine these e-processes by adjusting each and then averaging them.
@@
This is a challenging problem: Henzi and Law, (2024) show that it is impossible to construct nontrivial test martingales in the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0},\,\mathcal{F}_{t}=\sigma(\{(X_{i},Y_{i})\}_{i=1}^{t})$. (In fact, Ramdas et al., (2022) show the analogous statement for randomness testing.) For independence testing, the added challenge is that all known nontrivial e-processes, not just test martingales, are constructed by coarsening $\mathbb{F}$. Nevertheless, we still want to leverage the statistical power of different e-processes, so that we can detect different deviations from independence.
Recent work has developed two different types of e-processes for the independence null. First, Podkopaev et al., (2023) construct an e-process $\mathfrak{e}^{\mathsf{pair}}=(\mathfrak{e}_{t}^{\mathsf{pair}})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{indep}}$ based on the idea of betting on pairs of incoming data points (“pairwise betting”): $\mathfrak{e}_{0}^{\mathsf{pair}}=1$ and
@@
where $\lambda_{i}\in[-1,1]$ is ($\mathcal{F}_{2i-2}$)-measurable and $f_{i}\in[-1,1]$ is a payoff function that takes a pair of data points as inputs. The payoff is expected to be zero whenever $X_{2i}$ and $Y_{2i}$, as well as $X_{2i-1}$ and $Y_{2i-1}$, are independent. Crucial to our discussion is that this e-process is adapted in a sub-filtration that “updates” once every even number of steps: $\mathbb{F}^{[2]}\subseteq\mathbb{F}$, where $\mathbb{F}^{[k]}=(\mathcal{F}_{t}^{[k]})_{t\geq 0}\subseteq\mathbb{F}$ and $\mathcal{F}_{t}^{[k]}=\mathcal{F}_{\left\lfloor t/k\right\rfloor k}$. With this e-process, one can only stop at even times and not at odd times.
On the other hand, Henzi and Law, (2024) develop a test martingale $\mathfrak{e}^{\mathsf{rank}}=(\mathfrak{e}_{t}^{\mathsf{rank}})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{indep}}$ that is based on the sequential ranks of the data. Naturally, this e-process is only valid in a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$ generated by the ranks. In their experiments comparing the rank-based test martingale with the pairwise betting e-process, the authors remark that “neither of the methods uniformly dominate each other.” This motivates the use of a combined e-process.
@@
We note in closing that the recent work by Saha and Ramdas, (2024) also utilize pairwise betting but for testing randomness. The procedure described in this subsection would provide a way to combine their e-process with the UI and conformal e-processes in Section 4.
@@ 5.2 Evaluating and comparing $h$-step-ahead sequential forecasters
The next example represents a scenario where different e-processes, say $\mathfrak{e}^{[k]}$ for each $k$, are constructed in different sub-filtrations of the data filtration, say $\mathbb{F}^{[k]}\subsetneq\mathbb{F}$. In particular, we consider the problem of forecast evaluation, which is a classical task in econometrics, meteorology, and other areas (see, e.g., Diebold and Mariano,, 1995). Recent works have found useful synergies between this classical problem and the modern tools of SAVI. These works develop e-processes and sequential tests for testing whether one forecaster uniformly outperforms the other over time (Henzi and Ziegel,, 2022); testing whether a forecaster is calibrated (Arnold et al.,, 2023); and continuously monitoring the mean score difference between two forecasters (Choe and Ramdas,, 2023). It turns out that all three results implicitly utilize the $\mathfrak{p}$-lifting procedure; here, we build upon their results and describe how we can obtain e-processes directly via $\mathfrak{e}$-lifting.
We proceed with a concrete example based on the formulations of Choe and Ramdas, (2023, appendix E) and Arnold et al., (2023, section 4). Let $h>1$ be an integer lag, and suppose that at each time step $t=1,2,\dotsc$, two forecasters $\mathsf{p}$ and $\mathsf{q}$ make respective probability forecasts $p_{t},q_{t}\in[0,1]$ for an eventual binary outcome $y_{t+h-1}\in\{0,1\}$, say, whether it will rain on the (t+h1)(t+h-1)th day. The underlying distribution for the outcome sequence is not known. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration with which both $p_{t}$ and $q_{t}$ are $\mathcal{F}_{t-1}$-measurable—that is, $\mathcal{F}_{t-1}$ represents the information available to each forecaster before making their forecast on $y_{t+h-1}$.
@@
For example, if $h=2$, then the null asserts that $\mathsf{p}$ is no better than $\mathsf{q}$ on odd and even days. For each offset $k$, Choe and Ramdas, (2023) construct an e-process $\mathfrak{e}^{[k]}=(\mathfrak{e}_{t}^{[k]})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{pw}}$ in a sub-filtration $\mathbb{F}^{[k]}\subseteq\mathbb{F}$. The sub-filtration $\mathbb{F}^{[k]}$ updates its information every $h$ steps, starting at time $k$ (if $h=2$, then $\mathbb{F}^{[1]}=(\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{1},\mathcal{F}_{3},\mathcal{F}_{3},\mathcal{F}_{5},\mathcal{F}_{5},\dotsc)$). For $h>1$, this filtration design due to Arnold et al., (2023) is necessary because, at the time of forecasting $y_{t+h-1}$, we do not have access to the intermediary outcomes $y_{t},\dotsc,y_{t+h-2}$ that can affect any $\mathbb{F}$-stopping time.
@@
The $h$-step-ahead forecast comparison problem presents a unique application of $\mathfrak{e}$-lifting, as there are ways to bypass the construction of a combined e-process if one only cares about constructing a sequential test. In particular, as noted (implicitly) in prior work, we may directly utilize the “free” $\mathfrak{p}$-lifting procedure (Theorem 1) given the particular problem setup. To be clear, if our goal is to construct a combined e-process instead, then using an adjuster (25) remains the most straightforward and effective way. We elaborate on these points and include an empirical comparison of the resulting approaches, in Section H.2.
@@ 6 Characterizing adjusters as $\mathfrak{e}$-lifters
In Dawid et al., 2011a ; Dawid et al., 2011b ; Shafer et al., (2011); Koolen and Vovk, (2014), an adjuster is defined as a deterministic function that turns the running maximum of a test supermartingale into a process that is upper-bounded by another test supermartingale. These early works show that their definition is equivalent to the analytic definition of adjusters (11). We shall refer to their definition as game-theoretic because it views a test supermartingale as the wealth of a skeptic who bets against the null. In this view, an adjuster is a function that turns the running maximum of a skeptic’s wealth into a valid bet (i.e., the skeptic is not expected to make money if the data comes from the null). We include a game-theoretic interpretation of $\mathfrak{e}$-lifting in Section I.
Here, our main question is how this game-theoretic characterization of adjusters relates, if at all, to their ability to lift e-processes across filtrations. The first point of connection is that the existence of an upper-bounding test supermartingale is a characterizing property of e-processes (Ramdas et al.,, 2020, lemma 6). Then, in our context, an adjuster can be understood as a function that maps the running maximum of a test supermartingale, for $P$ in $\mathbb{G}$, into an e-process for $P$ in $\mathbb{G}$ (the same filtration). Then, by generalizing this definition to a condition that only involves e-processes, we will ultimately show how the game-theoretic notion is equivalent to a function being an $\mathfrak{e}$-lifter that can lift arbitrary e-processes across filtrations.

### v5

@@ 4 Experiments: Sequentially testing randomness
Given the $\mathfrak{e}$-lifting result (Theorem 2), we can now revisit the task of sequentially testing randomness, for which the issue of combining e-processes across filtrations arises. In this section, we first revisit our motivating example (Example 1) and show results on simulated data. Then, we apply our approach to testing the randomness of real-world financial time series data.
4.1 Simulations under the null and alternative hypotheses
@@
Recall from Example 1 that we could not simply average the two e-processes $\mathfrak{e}^{\mathsf{UI}}$ and $\mathfrak{e}^{\mathsf{conf}}$ for testing randomness, as the conformal test martingale (CTM) $\mathfrak{e}^{\mathsf{conf}}$ is not $\mathbb{F}$-anytime-valid. By Corollary 2, we can now combine the two e-process by lifting $\mathfrak{e}^{\mathsf{conf}}$ into $\mathbb{F}$ using an adjuster.
@@
As shown in the plots, unlike the original CTM, its adjusted version is mostly no greater than 1. Over $10,000$ repeated runs, the mean of the stopped e-values for the adjusted e-process is $0.60\pm 0.001$, suggesting that the adjusted e-process meets the validity definition (1) at this $\mathbb{F}$-stopping time. This validates our theoretical result that the adjusted e-process can be combined with the UI e-process via averaging, such that the combined version is an e-process in $\mathbb{F}$.
@@
Next, we generate data from a drifting distribution that is typically an IID $\mathsf{Ber}(0.5)$ (“usual state”), but after the first 500 steps it briefly switches to an IID $\mathsf{Ber}(0.2)$ (“unusual state”) for 100 steps, before returning to the usual state for the remainder of the sequence. The resulting data contains two changepoints over the sequence of length $T=2,500$. As mentioned earlier, CTMs are known to be powerful against changepoint alternatives, unlike the UI e-process. Figure 5(b), which shows the three e-processes averaged over 1,000 repeated runs, confirms this: the CTM grows large upon entering the unusual state and then gradually decreases upon reentering the usual state. In contrast, the UI e-process remains small throughout the sequence, unable to detect the brief state change. Because one of the two processes grows large (in the unusual state), analogous to the Markovian case, the combined e-process also grows large.
Unlike in Figure 5(a), however, in Figure 5(b) we additionally observe two peculiarities stemming from the $\mathfrak{e}$-lifting procedure. First, because the combined e-process utilizes the adjusted CTM, it grows less quickly ($\approx 10^{2}$ at $t=500$) than the CTM ($\approx 10^{4}$). In this example, the adjusted evidence still exceeds a typical threshold for a “strong” evidence (Shafer, 2021). (In a level-$\alpha$ sequential test (7) using e-processes, we would reject the null at $\alpha=0.05$ when the value of the e-process exceeds $20$, for example.) Second, as a byproduct of using an adjuster, the combined e-process retains the largest observed evidence over time (i.e., the running maximum). When the data returns to its usual state (past time 600), the base CTM gradually decreases, while the combined e-process eventually exceeds the CTM. Thus, while adjusting the e-process involves sacrificing some amount of evidence, it also comes with the useful byproduct of being able to insure against losing evidence over time.
@@ 4.2 Testing the randomness of high-volatility days in financial returns
To test the viability of the adjust-then-combine approach in real-world applications, we revisit and expand upon our motivating example (Example 1) involving financial time series data. As briefly described earlier, modeling the volatility of returns is a key problem in financial modeling, and a good volatility model is instrumental in tasks such as portfolio management, risk management, and options pricing (Engle and Patton, 2007). When considering a volatility model, it is useful to accompany it with a monitoring procedure that can check this model’s viability on live data, which may change unpredictably over time.
The problem of sequentially testing whether high-volatility days are distributed at random, as described in Example 1, is a concrete example of this monitoring task. A “stylized fact” about financial returns is that they are heteroskedastic, and specifically, that high-volatility days are often clustered together (e.g., Engle, 1982). This makes it practically important to detect exactly when (and how) volatility clustering happens on live data. The task can be viewed as a diagnostic test for whether a volatility model is needed in the first place: if high-volatility days were IID, then we would not expect a volatility model to provide added predictive power.
@@
In this experiment, we choose $c$ as the 80% upper quantile of the historical returns for the same asset, from a 1-year period before the first date (say, during the year 2018). In Figure 6, we plot the volatility process $V_{t}$ and the high-volatility days $X_{t}$ (marked as red) for the S&P 500 index, which shows volatility clustering around March 2020 and in 2022.
@@
While we focused on a binarized notion of volatility for simplicity, we can apply an analogous procedure to the continuous-valued residuals from a volatility model, by replacing the UI e-process with Saha and Ramdas (2024)’s e-process (powerful against autoregressive alternatives) and using the continuous analog of the CTM (Vovk, 2021b).
@@ 5.1 Combining e-processes for sequential independence testing
The first example is sequential (nonparametric) independence testing (Balsubramani and Ramdas, 2016; Podkopaev et al., 2023; Podkopaev and Ramdas, 2023; Shekhar and Ramdas, 2024; Henzi and Law, 2024). This represents a scenario where both e-processes are constructed in a coarsening of the data filtration. By a complete analogy to Corollary 2, we can combine these e-processes by adjusting each and then averaging them.
@@
This is a challenging sequential inference problem—Henzi and Law (2024) show that it is impossible to construct nontrivial test martingales in the data filtration $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0},\,\mathcal{F}_{t}=\sigma(\{(X_{i},Y_{i})\}_{i=1}^{t})$. (In fact, Ramdas et al. (2022) show the analogous statement for randomness testing.) Here, the added challenge is that all known nontrivial e-processes, not just test martingales, are constructed by coarsening $\mathbb{F}$. Nevertheless, we still want to leverage the statistical power of different e-processes, so that we can detect different deviations from independence.
Recent work has developed two different types of e-processes for the independence null. First, Podkopaev et al. (2023) construct an e-process $\mathfrak{e}^{\mathsf{pair}}=(\mathfrak{e}_{t}^{\mathsf{pair}})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{indep}}$ based on the idea of betting on pairs of incoming data points (“pairwise betting”): $\mathfrak{e}_{0}^{\mathsf{pair}}=1$ and
@@
where $\lambda_{i}\in[-1,1]$ is ($\mathcal{F}_{2i-2}$)-measurable and $f_{i}\in[-1,1]$ is a payoff function that takes a pair of data points as inputs. The payoff is expected to be zero whenever $X_{2i}$ and $Y_{2i}$, as well as $X_{2i-1}$ and $Y_{2i-1}$, are independent. Crucial to our discussion is that this e-process is adapted in a sub-filtration that “updates” once every even number of steps: $\mathbb{F}^{[2]}\subsetneq\mathbb{F}$, where $\mathbb{F}^{[k]}=(\mathcal{F}_{t}^{[k]})_{t\geq 0}$ and $\mathcal{F}_{t}^{[k]}=\mathcal{F}_{\left\lfloor t/k\right\rfloor k}$. With this e-process, one can only stop at even times and not at odd times.
On the other hand, Henzi and Law (2024) develop a test martingale $\mathfrak{e}^{\mathsf{rank}}=(\mathfrak{e}_{t}^{\mathsf{rank}})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{indep}}$ that is based on the sequential ranks of the data. Naturally, this e-process is only valid in a sub-filtration $\mathbb{G}\subseteq\mathbb{F}$ generated by the ranks. In their experiments comparing the rank-based test martingale with the pairwise betting e-process, the authors remark that “[…] neither of the methods uniformly dominate each other.” This motivates the use of a combined e-process.
@@
We note in closing that the recent work by Saha and Ramdas (2024) also utilize pairwise betting but for testing randomness. The procedure described in this subsection would provide a way to combine their e-process with the UI and conformal e-processes in Section 4.
@@ 5.2 Evaluating and comparing $h$-step-ahead sequential forecasters
The next example represents a scenario where different e-processes, say $\mathfrak{e}^{[k]}$ for each $k$, are constructed in different sub-filtrations of the data filtration, say $\mathbb{F}^{[k]}\subsetneq\mathbb{F}$. In particular, we consider the problem of forecast evaluation, which is a classical task in econometrics, meteorology, and other areas (see, e.g., Diebold and Mariano, 1995). Recent works have found useful synergies between this classical problem and the modern tools of SAVI. These works develop e-processes and sequential tests for testing whether one forecaster uniformly outperforms the other over time (Henzi and Ziegel, 2022); testing whether a forecaster is calibrated (Arnold et al., 2023); and continuously monitoring the mean score difference between two forecasters (Choe and Ramdas, 2023). It turns out that all three results implicitly utilize the $\mathfrak{p}$-lifting procedure; here, we build upon their results and describe how we can obtain e-processes directly via $\mathfrak{e}$-lifting.
We proceed with a concrete example based on the formulations of Choe and Ramdas (2023, appendix E) and Arnold et al. (2023, section 4). Let $h>1$ be an integer lag, and suppose that at each time step $t=1,2,\dotsc$, two forecasters $\mathsf{p}$ and $\mathsf{q}$ make respective probability forecasts $p_{t},q_{t}\in[0,1]$ for an eventual binary outcome $y_{t+h-1}\in\{0,1\}$, say, whether it will rain on the (t+h1)(t+h-1)th day. The underlying distribution for the outcome sequence is not known. Let $\mathbb{F}=(\mathcal{F}_{t})_{t\geq 0}$ be the filtration with which both $p_{t}$ and $q_{t}$ are $\mathcal{F}_{t-1}$-measurable—that is, $\mathcal{F}_{t-1}$ represents the information available to each forecaster before making their forecast on $y_{t+h-1}$.
@@
For example, if $h=2$, then the null asserts that $\mathsf{p}$ is no better than $\mathsf{q}$ on odd and even days. For each offset $k$, Choe and Ramdas (2023) construct an e-process $\mathfrak{e}^{[k]}=(\mathfrak{e}_{t}^{[k]})_{t\geq 0}$ for $\mathcal{H}_{0}^{\mathsf{pw}}$ in a sub-filtration $\mathbb{F}^{[k]}\subseteq\mathbb{F}$. The sub-filtration $\mathbb{F}^{[k]}$ updates its information every $h$ steps, starting at time $k$ (if $h=2$, then $\mathbb{F}^{[1]}=(\mathcal{F}_{0},\mathcal{F}_{1},\mathcal{F}_{1},\mathcal{F}_{3},\mathcal{F}_{3},\mathcal{F}_{5},\mathcal{F}_{5},\dotsc)$). For $h>1$, this filtration design due to Arnold et al. (2023) is necessary because, at the time of forecasting $y_{t+h-1}$, we do not have access to the intermediary outcomes $y_{t},\dotsc,y_{t+h-2}$ that can affect any $\mathbb{F}$-stopping time.
@@
The $h$-step-ahead forecast comparison problem presents a unique application of $\mathfrak{e}$-lifting, as there are ways to bypass the construction of a combined e-process if one only cares about constructing a sequential test. In particular, as noted (implicitly) in prior work, we may directly utilize the “free” $\mathfrak{p}$-lifting procedure (Theorem 1) given the particular problem setup. To be clear, if our goal is to construct a combined e-process instead, then using an adjuster (25) remains the most straightforward and effective way. We elaborate on these points and include an empirical comparison of the resulting approaches in Section H.2.
@@ 6 Characterizing adjusters as $\mathfrak{e}$-lifters
In Dawid et al. (2011a, b); Shafer et al. (2011); Koolen and Vovk (2014), an adjuster is defined as a deterministic function that turns the running maximum of a test supermartingale into a process that is upper-bounded by another test supermartingale. These early works show that their definition is equivalent to the analytic definition of adjusters (11). We shall refer to their definition as game-theoretic because it views a test supermartingale as the wealth of a skeptic who bets against the null. In this view, an adjuster is a function that turns the running maximum of a skeptic’s wealth into a valid bet (i.e., the skeptic is not expected to make money if the data comes from the null). We include a game-theoretic interpretation of $\mathfrak{e}$-lifting in Section I.
Here, our main question is how this game-theoretic characterization of adjusters relates, if at all, to their ability to lift e-processes across filtrations. The first point of connection is that the existence of an upper-bounding test supermartingale is a characterizing property of e-processes (Ramdas et al., 2020, lemma 6). Based on this characterization, an adjuster can be understood as a function that maps the running maximum of a test supermartingale, for $P$ in $\mathbb{G}$, into an e-process for $P$ in $\mathbb{G}$ (the same filtration). By generalizing the game-theoretic definition to a condition that only involves e-processes, we will show how the characterization is equivalent to a function being an $\mathfrak{e}$-lifter that can lift arbitrary e-processes across filtrations.